OpenStudy (anonymous):

specific solution for sin^2 theta+sin theta=0

OpenStudy (ipwnbunnies):

\[\sin^2 \theta + \sin \theta = 0\] \[\sin^2 \theta = -\sin \theta\] \[\sin \theta = -1\]